Suppose you know that Paul's company stock currently sells for $58 per share and the required return on the stock is 12.00%. You also know that the total return on the stock is evenly divided between capital gains yield and dividend yield. If its the company's policy to always maintain a constant growth rate in its dividends, what is the current dividend per share?
Answer #1

Dividend yield=Capital gains yield=(12/2)


=6%


Dividend yield=Dividend for next period/Current price


Dividend for next period=(58*6%)


=$3.48


Hence current dividend=Dividend for next period*Present value of discounting factor(rate%,time period)


=3.48/1.06


=$3.28(Approx)


answered by: Ace